A coffee merchant has coffee beans that sell for $9 a pound and $12 a pound. The two types are to be mixed to create 100 pounds of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
McGraw Hill 6th edition,8.2, # 60.
Thanky you for your time!!

Answer by stanbon(75887)

A coffee merchant has coffee beans that sell for $9 a pound and $12 a pound. The two types are to be mixed to create 100 pounds of a mixture that will sell for $11.25 per pound. How much of each type of bean should be used in the mixture?
--------------
Let amount of $9 coffee be "x" ; their value is 9x dollars
Amount of $12 coffee is "100-x" ; their value is 12(100-x)=1200-12x dollars
Amount of mixture is 100 lbs ; their value is 11.25(100)=1125 dollars
------------
EQUATION:
value + value = 1125 dollars
9x+1200-12x=1125
-3x=-75
x=25 (lbs. of $9 coffee)
100-x=75 (lbs. of $12 coffee)
-----------
Cheers,
Stan H.